# Towards Large-Scale In-Context Reinforcement Learning by Meta-Training in Randomized Worlds

### v2

@@ 3.2 Modeling and Training Framework
Figure 1: Comparison of the learning pipelines and data formulation of different ICRL methods (AD, ADε,DPT) and our methods (step-wise supervision, SS), all of which are based on supervised learning.
@@
Step-wise supervision (SS): Meta-training for ICRL using RL or Evolution Strategies faces significant challenges, including cumbersome infrastructure requirements and high computational costs. Recent advances in supervised learning-based meta-training methods—such as Algorithm Distillation (AD) [12], ADε [14], and Decision Pre-Training Transformers (DPT) [15]—have shown promise for scalable ICRL meta-training. However, a critical challenge arises during inference: contextual trajectories are generated by the model itself, creating an unavoidable gap between training-time and inference-time trajectories. This discrepancy can lead to catastrophic failures during deployment. While leveraging diverse behavior trajectories has been shown to mitigate this issue [15], it introduces a new challenge of maintaining training efficiency. To address these limitations, we propose step-wise supervision (SS), a framework inspired by DPT and data aggregation techniques for imitation learning [61, 62]. Our approach hinges on two key policies: The behavior policy refers to the policy executed to generate trajectories during training. The reference policy refers to the target policy to be imitated, which remains decoupled from direct execution. Decoupling the behavior and reference policies enables the introduction of diversity into the behavior policy, thereby reducing the discrepancy between training and inference trajectories while maintaining the optimality of the reference policy, as shown in Figure 1. Unlike DPT, which imitates only one-step action conditioned on a trajectory, our step-wise supervision framework is inherently designed to align with high-efficiency chunk-wise training pipelines for sequence models such as Transformers [63] and their optimized variants [64, 65].
@@
Inducing prior knowledge into ICRL: For SS, which employs diverse policies for trajectory generation, prior knowledge becomes crucial for interpreting actions derived from heterogeneous policies [66]. This motivates the incorporation of prior knowledge, specifically metadata indicating the policy used to generate each action. In this work, we implement a diverse set of behavior policies $\Pi=\{\pi^{(b)}\}$, with $(b)$ denoting the behavior policies. $\pi^{(b)}$ comprises seven distinct types, including myopic greedy, oracle, Q-learning, and model-based reinforcement learning, denoted by a marker $tag(\pi)\in\{0,...,6\}$. To handle unseen or unclassified policies, we reserve an additional identifier “Unk” with $m=7$. The trajectory is denoted by $h_{T}(\tau,\pi)=[(s_{1},g_{1},a_{1},r_{1}),...,(s_{t},g_{T},a_{T},r_{T})]$, where $s_{t}\in\mathcal{S}$, $a_{t}\in\mathcal{A},a_{t}\sim\pi^{(b)}(s_{t})$, $r_{t}\sim\mathcal{R}(s_{t},a_{t},s_{t+1})$, and $g_{t}=g(a_{t})=tag(\pi^{(b)})$ denotes the prior knowledge for action.
@@
Data synthesis and meta-training: We use $\pi*$ to denote the reference policy, which is the oracle policy with $\gamma>0.99$. It is then used to label a list of the reference actions step-by-step as $l_{T}(\tau)=[a^{*}_{1},a^{*}_{2},...,a^{*}_{T}]$ with $a^{*}_{t}\sim p^{*}_{t}=\pi^{*}(s_{t})$. Notice that for most of the time $a_{t}\neq a^{*}_{t}$. We first collect the training and validation datasets by:
@@
We then meta-train the model in the dataset by optimizing the following target:
$p^{\theta}_{1},p^{\theta}_{2},...,p^{\theta}_{t}$ $=Causal_{\theta}(s_{0},g_{0},a_{0},r_{0},s_{1},...,s_{2},...,...,s_{t}).$ (3)
@@
Figure 2: OmniRL model structure and training framework
which is also shown in Figure 2.
Scaling up the trajectory length: Extending ICL to complex tasks at scale requires efficient modeling of very long contexts. While the Transformer [63] suffers from both memory limits and computation cost limits for long sequences, we leverage linear attention architectures, such as gated slot attention (GSA) [65] layers and RWKV-7 [67], to extend sequence length scalability, achieving linear computational scaling during inference. To further break down the limitation in context length, we break a long sequence $h_{T}$ into K segments $[0,T_{1}],[T_{1}+1,T_{2}],...,[T_{K-1}+1,T_{K}]$. The forward pass is calculated recurrently across the segments, and the backward calculation is performed within each segment. The gradient for the memory states of the linear attention layer $\phi_{t}$ is blocked across the segments. Equation 3 is thereby replaced by:
with $SG$ representing stopping gradient. In the meta-training process, the gradients are calculated within each segment and accumulated in cache first. They are applied to the parameters only at the end of the trajectory.
@@ 4.1 Demonstration of Generalization and Scalability
Table 1: We pick the result achieving its best episode performance within 10,000 episodes for each model, and list the performances, the average steps and episodes required to achieve the score. The hyper-parameters for Q-Learning and PPO are optimized under the evaluated task or task set. The episode performances are normalized to a scale of $0.0$ (uniform random policy) to $1.0$ (oracle policy). Results for AnyMDP tasks $\mathcal{T}_{tst}$ are averaged over $64$ tasks per class, while results for Gymnasium tasks are averaged over $3$ independent runs for each task. Notably, all evaluated tasks do not overlap with the OmniRL training task set; furthermore, Gymnasium, DarkRoom, and Bandits tasks are entirely absent from OmniRL’s training regimen.
@@
TQL-UCB PPO OmniRL(Ours)
$\mathcal{T}_{tst}(1,5)$ (Bandits) $92.1\%/100/100$ $95.6\%/1.2K/1.2K$ $82.5\%/103/103$
$\mathcal{T}_{tst}(16,5)$ $92.0\%/297K/4.7K$ $90.6\%/476K/9.7K$ $2995.3\%/2.0K/$ 29
$\mathcal{T}_{tst}(32,5)$ $84.7\%/616K/5.6K$ $72.2\%/618K/9.7K$ $4790.3\%/6.5K/$ 47
$\mathcal{T}_{tst}(64,5)$ $83.7\%/1.1M/5.1K$ $58.3\%/1.1M/9.4K$ $2591.3\%/7.7K/$ 25
$\mathcal{T}_{tst}(128,5)$ $73.2\%/1.8M/6.9K$ $49.0\%/1.3M/8.6K$ $10080.2\%/36.3K/$ 100
CliffWalking $35100\%/3.1K/$ 35 $95.9\%/99.3K/2.7K$ $65100\%/3.0K/$ 65
FrozenLake (non-slippery) $95.3\%/23.6K/3.7K$ $96.8\%/18.2K/2.1K$ $3599.8\%/0.3K/$ 35
FrozenLake (slippery) $96\%/208K/10.0K$ $95.6\%/73.6K/4.7K$ $24579.5\%/7.7K/$ 245
Discrete-Pendulum (g=1) $11094.9\%/22K/$ 110 $99099.3\%/198K/$ 990 $4090.5\%/8K/$ 40
Discrete-Pendulum (g=5) $99.7\%/426K/2.13K$ $66099.8\%/132K/$ 660 $17091.8\%/34K/$ 170
Discrete-Pendulum (g=9.8) $90.2\%/2.0M/10.0K$ $93098.3\%/186K/$ 930 $16573.4\%/33K/$ 165
Switch2 (Multi-Agent)[68] $11098\%/3.8K/$ 110 $10080.4\%/2.8K/$ 100
Darkroom (6x6) $48198.1\%/6.2K/$ 481 $56097.6\%/10.6K/$ 560 $95.2\%/845/40$
Darkroom (8x8) $96.8\%/24.5K/2.0K$ $93096.7\%/15.9K/$ 930 $4093.8\%/1.5K/$ 40
Darkroom (10x10) $89\%/31.1K/1.7K$ $57092.3\%/15.7K/$ 570 $10091.7\%/2.8K/$ 100
Figure 3: Performance comparison of TQL-UCB, PPO, and OmniRL on test tasks $|\mathcal{T}_{tst}|=64$. The shaded regions denote $95\%$ confidence intervals. Online evaluation performance is measured across episodes. Grid search on the hyperparameters for PPO and TQL-UCB is conducted independently within each task group to achieve their best performance at the final episodes. Additionally, we plot the step-wise validation loss of OmniRL on validation dataset $\mathcal{D}(\mathcal{T}_{tst})$, generated using the same data synthesis methodology as the training dataset but applied to test tasks. The validation loss exhibits strong alignment with online-RL evaluation results.
@@
Figure 4: Comparison of meta-training dynamics across AnyMDP dataset of $6B$ time steps and step-wise validation loss for different model architectures.
@@
The meta-training process is primarily conducted using 8 Nvidia Tesla A800 GPUs. We use a batch size of $5$ per GPU, divided into segments (chunks) of 2K steps each. We optimize using the AdamW algorithm with a learning rate that decays from a peak value of $2\times 10^{-4}$. The average time cost per iteration is $8$ seconds for trajectories with $T=12K$, and this cost increases linearly with sequence length. For more details please check Section C.2. For the causal sequence model, we evaluate four architectures: RWKV-7 [67], Gated Delta-Net (GDN) [69], Gated Self-Attention (GSA) [65], Mamba2 [70]. In previous tests, model architectures of the Linear Attention type outperformed Transformers in both training efficiency and context memory length. Therefore, the experiments in the figure Figure 5 focus exclusively on linear-attention architectures. The test results are largely consistent with the conclusions in language processing, demonstrating the capability of AnyMDP to serve as a benchmark for long-term sequence modeling. After comparing computational efficiency and performance, we select RWKV-7 for subsequent experiments.
@@
Without any further parameter tuning, we evaluate our model, namely OmniRL, on both unseen AnyMDP tasks in Figure 3, Gymnasium tasks, and DarkRoom [12] in Figure 16, and those performances are shown in Table 1. Notably, unlike previous ICRL works, our training set does not include any instances of DarkRoom. In our experiments, the selected tasks are constrained to environments with observation spaces of dimension $n_{s}\leq 128$ and action spaces of dimension $n_{a}\leq 5$. For environments with continuous observation spaces, such as Pendulum-v1, we manually discretize the observation space into 60 discrete classes using a grid-based discretization method. To adapt OmniRL that is trained with $n_{a}=5$ to environments with less actions ($n_{a}<5$), we reassign unused actions to valid ones. This further demonstrates the compatibility of OmniRL across environments with varying action space dimensions. We also found that proper reward shaping is important for OmniRL to work, as shown in Figure 14; the details can be found at Section C.3.
@@
In Table 1, we compare the normalized performance, episode cost, and step costs of OmniRL, classical Tabular Q-learning (TQL) [71] with upper confidence bound (UCB) [72] (TQL-UCB for short), and Proximal Policy Optimization (PPO) [73]. Results show that OmniRL, when meta-trained exclusively on the proposed AnyMDP tasks, effectively adapts to most Gymnasium tasks, validating the representational capability of AnyMDP tasks. This result also demonstrates OmniRL’s superior sample efficiency, which aligns with prior ICRL findings. Notably, despite being trained solely on single-agent tasks, OmniRL adapts to multi-agent tasks like Switch2 by configuring observation spaces, enabling emergent inter-agent cooperation without explicit multi-agent interaction during training and thus decoupling cooperative behavior emergence from centralized mechanisms. Furthermore, in line with expectations, solving AnyMDP tasks becomes more difficult with increased state($n_{s}$) or action($n_{a}$) space size, with PPO proving more sensitive to action space extension and TQL-UCB more vulnerable to state space growth, as illustrated in Figure 9.
Visualizing ICRL. ICRL with linear attention captures all the information required to solve the environment in its memories ($\phi_{t}$). We perform a comprehensive t-SNE analysis to examine how these memories transform across different tasks during Online-RL evaluation. As shown in Figure 5, the clustering patterns confirm the distinct task distributions of Gym, Darkroom, and AnyMDP. Notably, Darkroom and Gym clusters are predominantly located in the top-left region, while AnyMDP occupies a broader spatial area, reflecting its greater diversity. This spatial differentiation emphasizes AnyMDP’s unique characteristics and highlights OmniRL’s strong generalization ability across diverse tasks.
@@ 4.2 OmniRL Performs Both Offline and Online Learning Better
For the ablation study and comparison with the other methods including AD, ADϵ, and DPT, we collect a smaller dataset with $|\mathcal{D}_{small}|$ comprising $128K$ sequences for training, where $n_{s}=16,n_{a}=5,T=8K$, with a total of $1B$ time steps. Figure 6 summarizes the performance of different methods trained on $\mathcal{D}_{Small}$ with identical training iterations. The comparison includes AD, ADε, DPT, OmniRL, and OmniRL (w/o a priori) where the prior info $g_{t}$ is removed from the sequence.
We examine the performance of different methods with different initial contexts: (1) Online-RL: The agent starts with an empty trajectory ($h_{0}=\emptyset$). (2) Offline-RL: The agent starts with an existing context derived from imperfect demonstrations (e.g., disturbed oracle policy) ($h_{0}=h^{\pi}$). (3) Imitation Learning: The agent starts with an existing context derived from oracles($h_{0}=h^{(exp)}$). For all three categories, the subsequent interactions are continually added to the context. Therefore, the models differ only in their initial memory or cache. The evaluation assesses the agents’ abilities in two key areas: their capacity to exploit existing information and their ability to explore and exploit continually. In the results in Figure 6, OmniRL and OmniRL (w/o a priori) surpass AD, ADε, and DPT with large gap, validating the effectiveness of Step-wise Supervision (SS). OmniRL (w/o a priori) lags behind OmniRL with a noticeable gap in all three groups, demonstrating the effectiveness of integrating the prior information. Table 2 and Figure 17 further demonstrate that the offline-learning ability of OmniRL can generalize to Gymnasium environments.
Figure 6: Evaluation results of AD, ADϵ, DPT, and OmniRL on 32 AnyMDP tasks, with 3 groups of initial demonstrations to assess the capabilities of online-RL, offline-RL, and imitation learning. For offline-RL and imitation learning, the agent is initialized with the context including $100$ episodes of demonstration. The results demonstrate the effectiveness of Step-wise Supervision (SS) and the integration of prior information.
Table 2: We select four Gymnasium environments to evaluate the performance of Offline RL and Imitation Learning. In each environment, both oracle and random teacher demonstrations are provided. Conservative Q-Learning (CQL) [74] updates its value table using the demonstration data, while OmniRL loads the data in its memory. The performance across the first 0–20 episodes reflects the effectiveness of Offline RL and Imitation Learning, whereas the performance in episodes 180–200 demonstrates the outcomes after online learning.
Environments Teacher(Performance) 0$\sim$20/180$\sim$200 Episodes Performance

### v3

@@ 4 The Scalable ICRL Framework of OmniRL
Figure 3: Comparison of the learning pipelines and data formulation of different ICRL methods (AD, ADε,DPT) and Decoupled Policy Distillation (DPD).
@@
Decoupled Policy Distillation (DPD): Meta-training for ICRL using RL or Evolution Strategies faces significant costs, including cumbersome infrastructure requirements and high computational costs. Recent advances in supervised learning-based meta-training methods—such as Algorithm Distillation (AD) [13], ADε [15], and Decision Pre-Training Transformers (DPT) [16]—have shown promise for scalable ICRL meta-training. However, a critical challenge arises during inference: contextual trajectories are generated by the model itself, creating an unavoidable gap between training-time and inference-time trajectories. This discrepancy can lead to catastrophic failures during deployment. While leveraging diverse behavior trajectories has been shown to mitigate this issue [16], it introduces a new challenge of maintaining training efficiency. To address these limitations, we propose Decoupled Policy Distillation (DPD), a framework inspired by DPT and data aggregation techniques for imitation learning [62, 63]. Our approach hinges on two decoupled policies: The behavior policy refers to the policy executed to generate trajectories during training. The reference policy refers to the target policy to be imitated, which remains decoupled from direct execution. Decoupling the behavior and reference policies enables the introduction of diversity into the behavior policy, thereby reducing the discrepancy between training and inference trajectories while maintaining the optimality of the reference policy, as shown in Figure˜3. Unlike DPT, which imitates only one-step action conditioned on a trajectory, our step-wise supervision framework is inherently designed to align with high-efficiency chunk-wise training pipelines for sequence models such as Transformers [64] and their optimized variants [65, 66], producing $T$ times training efficiency ($T$ is the sequence length).
@@
Prior knowledge augmented ICRL: For SS, which employs diverse policies for trajectory generation, prior knowledge becomes crucial for interpreting actions derived from heterogeneous policies [67]. This motivates the incorporation of prior knowledge, specifically metadata indicating the policy used to generate each action. In this work, we implement a diverse set of behavior policies $\Pi=\{\pi^{(b)}\}$, with $(b)$ denoting the behavior policies. $\pi^{(b)}$ comprises seven distinct types, including myopic greedy, oracle, Q-learning, and model-based reinforcement learning, denoted by a marker $tag(\pi)\in\{0,...,6\}$. To handle unseen or unclassified policies, we reserve an additional identifier “Unk” with $m=7$. The trajectory is denoted by $h_{T}(\tau,\pi)=[(s_{1},g_{1},a_{1},r_{1}),...,(s_{t},g_{T},a_{T},r_{T})]$, where $s_{t}\in\mathcal{S}$, $a_{t}\in\mathcal{A},a_{t}\sim\pi^{(b)}(s_{t})$, $r_{t}\sim\mathcal{R}(s_{t},a_{t},s_{t+1})$, and $g_{t}=g(a_{t})=tag(\pi^{(b)})$ denotes the prior knowledge for action.
@@
Chunwise Training: We use $\pi^{*}$ to denote the reference policy, which is the oracle policy with $\gamma>0.99$. It is then used to label a list of the reference actions step-by-step as $l_{T}(\tau)=[a^{*}_{1},a^{*}_{2},...,a^{*}_{T}]$ with $a^{*}_{t}\sim p^{*}_{t}=\pi^{*}(s_{t})$. Notice that for most of the time $a_{t}\neq a^{*}_{t}$. We first collect the training and validation datasets by:
@@
Scaling ICL to handle complex tasks necessitates the efficient modeling of extensive contexts, which in turn demands substantial computational resources and large-capacity hardware memories. To further break down the limitation in context length, we break a long sequence $h_{T}$ into K segments $1,T_{1}],[T_{1}+1,T_{2}],...,[T_{K-1}+1,T_{K}]$. The forward pass is calculated recurrently across the segments, and the backward calculation is performed within each segment. The gradient for the memory states of the linear attention layer $\phi_{t}$ is blocked across the segments. Thus the sequence is encoded in the following chunkwise manner:
with $SG$ representing stopping gradient. In the meta-training process, the gradients are calculated within each segment and accumulated in cache first. They are applied to the parameters only at the end of the trajectory. The final target is as follows:
@@
The complete architecture of the model and its training pipeline are depicted in Figure˜4. We utilize a causal sequence model (where the prediction at each position relies solely on information from preceding positions) to encode the input sequence and forecast the action at the position aligned with the state inputs. In this paper, we mainly employ linear attention layers including gated slot attention (GSA) [66] layers, mamba-2 [68], and RWKV-7 [69] as the causal sequence models.
Figure 4: OmniRL model structure and training losses.
@@ 5.1 Demonstration of Generalization and Scalability
Table 1: Comparison of best average episodic performance within 10,000 episodes for each learner. The episode performances are normalized to a scale of $0\%$ (uniform random policy) to $100\%$ (oracle policy). The minimum steps and episodes required to achieve within at least $99\%$ of the best episodic performances are also listed. The hyper-parameters for Q-Learning and PPO are optimized under the evaluated task or task set. Results for AnyMDP and Garnet Tasks are averaged over $64$ tasks, results for Gymnasium tasks are averaged over 33 independent runs for each task. Gymnasium, DarkRoom, and Bandits tasks are entirely absent from OmniRL’s training regimen.
@@
TQL-UCB PPO
OmniRL
OmniRL
$\mathcal{T}_{tst}(1,5)$ (Bandits) $92.1\%/100/100$ $95.6\%/1.2K/1.2K$ $82.5\%/103/103$ $46.6\%/33/33$
$\mathcal{T}_{tst}(16,5)$ $92.0\%/297K/4.7K$ $90.6\%/476K/9.7K$ $95.3\%/2.0K/29$ $47.8\%/1.6K/24$
$\mathcal{T}_{tst}(32,5)$ $84.7\%/616K/5.6K$ $72.2\%/618K/9.7K$ $90.3\%/6.5K/47$ $42.0\%/5.0K/44$
$\mathcal{T}_{tst}(64,5)$ $83.7\%/1.1M/5.1K$ $58.3\%/1.1M/9.4K$ $91.3\%/7.7K/25$ $47.1\%/6.6K/24$
$\mathcal{T}_{tst}(128,5)$ $73.2\%/1.8M/6.9K$ $49.0\%/1.3M/8.6K$ $80.2\%/36.3K/100$ $32.3\%/9.0K/31$
Garnet(16,5,2)(16,5,2) $98.8\%/241K/2.1K$ $97.1\%/57K/0.5K$ $85.9\%/8.2K/71$ $99.0\%/10.8K/95$
Garnet(64,5,2)(64,5,2) $98.7\%/614K/1.7K$ $98.1\%/96K/0.26K$ $80.4\%/8.0K/19$ $87.3\%/7.4K/23$
CliffWalking $100\%/3.1K/35$ $95.9\%/99.3K/2.7K$ $100\%/3.0K/65$ $63\%/29K/300$
FrozenLake (non-slippery) $95.3\%/23.6K/3.7K$ $6.8\%/18.2K/2.1K 99$$9.8\%/0.3K/35$ $75.1\%/4.0K/250$
FrozenLake (slippery) $96\%/208K/10.0K$ $95.6\%/73.6K/4.7K$ $79.5\%/7.7K/245$ $31.3\%/11.8K/800$
Discrete-Pendulum (g=1) $4.9\%/22K/110 99$$9.3\%/198K/990$ $90.5\%/8K/40$ $0.0\%/-/$-
Discrete-Pendulum (g=5) $9.7\%/426K/2.13K 99$$9.8\%/132K/660$ $91.8\%/34K/170$ $0.0\%/-/$-
Discrete-Pendulum (g=9.8) $90.2\%/2.0M/10.0K$ $98.3\%/186K/930$ $73.4\%/33K/165$ $0.0\%/-/$-
Switch2 (Multi-Agent)[70] $98\%/3.8K/110$ $80.4\%/2.8K/100$
Darkroom (6x6) $98.1\%/6.2K/481$ $97.6\%/10.6K/560$ $95.2\%/845/40$ $90.5\%/21.3K/440$
Darkroom (8x8) $96.8\%/24.5K/2.0K$ $96.7\%/15.9K/930$ $93.8\%/1.5K/40$ $88.9\%/30.4K/480$
Darkroom (10x10) $89\%/31.1K/1.7K$ $92.3\%/15.7K/570$ $91.7\%/2.8K/100$ $75.6\%/20.8K/280$
Figure 5: Left: Normalized episodic return of TQL-UCB, PPO and OmniRL on 64 AnyMDP evaluation tasks (mean $\pm 95\%$ CI). Hyper-parameters of PPO and TQL-UCB were separately tuned by grid-search inside every task family to maximize final-episode performance. Right: Per-step validation loss of OmniRL on a held-out static dataset, mirroring the online-RL trend.
@@
The meta-training process is primarily conducted using 8 Nvidia Tesla A800 GPUs. We use a batch size of $5$ per GPU, divided into segments (chunks) of 2K steps each. We optimize using the AdamW algorithm with a learning rate that decays from a peak value of $2\times 10^{-4}$. The average time cost per iteration is $8$ seconds for trajectories with $T=12K$, and this cost increases linearly with sequence length. For more details please check Section˜C.2. For the causal sequence model, we evaluate four architectures: RWKV-7 [69], Gated Delta-Net (GDN) [71], Gated Self-Attention (GSA) [66], Mamba2 [72]. The test results are largely consistent with the conclusions reported in language processing (RWKV-7 $\approx$ GDN >> GSA >> Mamba2, with details in Section˜D.1), demonstrating the capability of AnyMDP to serve as a benchmark for long-term sequence modeling. Therefore, we select RWKV-7 for subsequent experiments.
@@
Without any further parameter tuning, we evaluate our model, namely OmniRL, on both unseen AnyMDP tasks in Figure˜5, Gymnasium tasks, and DarkRoom [13] in Figure˜16, and those performances are shown in Table˜1. Notably, unlike previous ICRL works, our training set does not include any instances of DarkRoom. In our experiments, the selected tasks are constrained to environments with observation spaces of dimension $n_{s}\leq 128$ and action spaces of dimension $n_{a}\leq 5$. For environments with continuous observation spaces, such as Pendulum-v1, we manually discretize the observation space into 60 discrete classes using a grid-based discretization method. To adapt OmniRL that is trained with $n_{a}=5$ to environments with less actions ($n_{a}<5$), we reassign unused actions to valid ones. This further demonstrates the compatibility of OmniRL across environments with varying action space dimensions. We also found that proper reward shaping is important for OmniRL to work, as shown in Figure˜13; the details can be found at Section˜C.3.
@@
In Table˜1, we compare the normalized performance, episode cost, and step costs of OmniRL, classical Tabular Q-learning (TQL) [59] with upper confidence bound (UCB) [60] (TQL-UCB for short), and Proximal Policy Optimization (PPO) [61]. OmniRL, meta-trained solely on AnyMDP, adapts effectively to most unseen tasks, confirming AnyMDP’s representational power; by contrast, OmniRL trained on Garnet MDPs excels only on Garnet MDPs. The results also demonstrates OmniRL’s superior sample efficiency, which aligns with prior ICRL findings. Notably, despite being trained solely on single-agent tasks, OmniRL adapts to multi-agent tasks like Switch2 by configuring observation spaces, enabling emergent inter-agent cooperation without explicit multi-agent interaction during training and thus decoupling cooperative behavior emergence from centralized mechanisms. Furthermore, in line with expectations, solving AnyMDP tasks becomes more difficult with increased state($n_{s}$) or action($n_{a}$) space size, with PPO proving more sensitive to action space extension and TQL-UCB more vulnerable to state space growth, as illustrated in Figure˜9.
@@ 5.2 OmniRL Performs Both Offline and Online Learning Better
For the ablation study and comparison with the other methods including AD, ADϵ, and DPT, we collect a smaller dataset with $|\mathcal{D}_{small}|$ comprising $128K$ sequences for training, where $n_{s}=16,n_{a}=5,T=8K$, with a total of $1B$ time steps. Figure˜6 summarizes the performance of different methods trained on $\mathcal{D}_{Small}$ with identical training iterations. The comparison includes AD, ADε, DPT, OmniRL, and OmniRL (w/o a priori) where the prior info $g_{t}$ is removed from the sequence.
[…]
Figure 6: AD, ADϵ, DPT, and OmniRL on 32 AnyMDP tasks, each tested under three initial contexts tailored for online RL, offline RL, and imitation learning. offline-RL and IL agents are seeded with 100 demonstration episodes before any interaction begins; results highlight the gains from DPD and prior-information integration.
Table 2: Offline-RL and IL performance of meta-trained OmniRL versus Conservative Q-Learning (CQL) [73] on four unseen Gymnasium tasks, using demonstrations from both oracle and random policies as initial contexts. Returns in episodes 0–20 measure offline RL and IL efficacy; returns in episodes 180–200 show the benefit of subsequent online RL.
[…]